OpenStudy (anonymous):

Match the functions (f) with there inverse (f^-1). f(x)= 3x + 1 f(x)=\frac{7 \cdot x-4}{8} f(x)=3x + 3 f(x)= \frac{1}{3} x + 4 f(x) = \frac{x+6}{2} Answer Choices 3x-12 2x-6 no inverse listed (x-3)/3 (8x+4)/7 (x-1)/3

OpenStudy (anonymous):

for each of these substitute x in for y so like the first one instead of y=3x+1 do x=3y+1 and resolve for y.

OpenStudy (anonymous):

and what would you get?

OpenStudy (anonymous):

I'm not sure...I don't understand how to do it

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok. good luck with the rest! if you have troubles just ask again.

OpenStudy (anonymous):

and i made one error in the previous post so i will post again: \[f(x)=3x+1\] \[y=3x+1\] \[x=3y+1\] \[x-1=3y\] \[\frac{ x-1 }{ 3 }=y\] \[f^{-1}(x)=\frac{ x-1 }{ 3 }\]

OpenStudy (anonymous):

yes it would be :)

OpenStudy (anonymous):

ahh.. ok I've got it know thank you

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

I've did the others but I can't find the answer to the second or last one

OpenStudy (anonymous):

\[f(x)=\frac{7 \cdot x-4}{8} \] is the second one so first switch the x and y \[x=\frac{7 \cdot y-4}{8} \] and multiply by 8 to get \[8x=7y-4\] Add 4 \[8x+4=7y\] and divide by 7 to get \[f^-1(x)=\frac{ 8x+4 }{ 7 }\]

OpenStudy (anonymous):

Now for the last one \[f(x) = \frac{x+6}{2} \] \[x=\frac{ y+6 }{ 2 }\] \[2x=y+6\] \[2x-6=y\] \[f^-1(x)=2x-6\] Hope this helps!

OpenStudy (anonymous):

thank you it does!